1. A plane is ruled with parallel straight lines at equal distances of 2a. A needle, 2L long(L



2. A lot contains 20 articles.The probablility that the lot contains exactly 2 defective articles is 0.4 and that the lot contains exactly 3 defective articles is 0.6 . Articles are drawn from the lot at random one by one without replacement and are tested till all the defective articles are found. What is the probability that the testing procedure ends at th 12th testing.



Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sudip

let

A=testing procedure ends at 12th testing

A1 =lot contain 2 defective article

A2 =lot contain 3 defective article

required probability=P(A)

                            =P(A∩ A1)∩ (A∩A2)

                            =P(A∩ A1)+(A∩A2)

                           =P(A1)P(A/A1)  + P(A2)P(A/A2)

 

now  P(A/A1)=probability that first 11 draws contain 10 non defective and 1 defective and 12th draw contain a defective article

                 =18C10 X2C1/20C11  X  1/9

 

 and P(A/A1)=probability that first 11 draws contain 9  non defective and 2 defective and 12th draw contain a defective article

                 =17C9 X3C2/20C11  X  1/9

Hence required probability

    = 0.4 X 18C10 X2C1/20C11  X  1/9        +0.6 X 17C9 X3C2/20C11  X  1/9

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin